Assume an isoquant for a fixed level of output equal to ¯y = 1/2x 1 2 1 x 1 3 2 Fix w1 = 1. Show how the cost minimizing bundle changes as w2 moves from 1 to 2. You can graph x1 and x2 on separate graphs as a function of w.\
Answer #1

hower Given that . = 4x4 ch.9 -0. g isoquant to,= ! The cost miniming barndo changes as as moren from 1902. -> ISO Quant isoc- I od da, : 24/24 - 24/2d2 2 -2.2, ½ a 113 622sail2. =61422. is principle of cost minimization - 301 r. Q2=-2/3. wal, w2=3.Cost minining bundle now had Cons a, and more ni rm The to need from stubstituta relatively cheapee input relatively costlier